(VW+WX)/VW=(VZ+YZ)/VZ what would this be called?
1
Expert's answer
2019-10-01T10:59:59-0400
"\frac{VW+WX}{VW}=\frac{VZ+YZ}{VZ}\ or\\\\"



"1+\frac{WX}{VW}=1+\frac{YZ}{VZ}"



"\frac{WX}{VW}=\frac{YZ}{VZ}"


So we can call that fraction of WX to VW is equal to that of YZ to VZ